The velocity of a solar powered car as a function of time. The driver accelerates from a stop sign. Cruises for 20s at a constant speed of 60km/hr and then brakes to come to a stop 40s after leaving the stop sign
a) Compare the average acceleration during the following time intervals
i) t=0 to t=10s;
ii) t= 30s to t=40s
iii) t=10 to t= 30s
iv) t=0 to t=40s
b) what is the instantaneous acceleration at t = 20s and t=35s ?
1
Expert's answer
2020-03-18T10:15:19-0400


a) i)


"a=\frac{60}{3.6\cdot 10}=1.67\frac{m}{s^2}"

ii)


"a=\frac{0-60}{3.6\cdot 10}=-1.67\frac{m}{s^2}"

iii)


"a=0\frac{m}{s^2}"

iv)


"a=\frac{0-0}{40}=0\frac{m}{s^2}"

b)


"a(20)=0\frac{m}{s^2}"

"a(35)=-1.67\frac{m}{s^2}"